OpenStudy (anonymous):

y=x^2+2x+1 find the vertex

OpenStudy (anonymous):

2x + 2 = 0 x = -1 (-1, 0)

OpenStudy (anonymous):

vertex is \[(-1,0)\] because \[y=x^2+2x+1=(x+1)^2\]

OpenStudy (anonymous):

in general vertex is \[x=-\frac{b}{2a}\] and then y is whatever you get when you replace x by \[-\frac{b}{2a}\]

OpenStudy (anonymous):

the axis of symmetry?

OpenStudy (anonymous):

it is \[x=-\frac{b}{2a}\] again. in this case it is \[x=-1\]

OpenStudy (anonymous):

would that be a minimum or maximum value? and what would the value be? and what is the range of the parabola

OpenStudy (anonymous):

these are all essentially the same question, just written in different forms. a) the vertex is (-1,0) b) the axis of symmetry is x = -1 c) the minimum value is 0 (the second coordinate of a vertex, and it is the minimum because this parabola opens up d) the range is \[[0,\infty)\] for the same reason as answer c

OpenStudy (anonymous):

thank you:)!

OpenStudy (anonymous):

thank you:)!

OpenStudy (anonymous):

yw